Consider an urn in which 4 balls have been placed by the following scheme : A fair coin
is tossed; if the coin falls head, a white ball is placed in the urn, and if the coin falls tail, a red
ball is placed in urn. (
i)What is the probability that the urn will contain exactly3 white balls ?
 

Grade:12th pass

1 Answers

Athul
6 years ago
Probability of No of white balls being 3 is same as probability of getting 3 heads ......To find the p (3 heads): n(3H)/n(S). = 4/16=1/4..... And:1/4